OpenStudy (princess123):

When simplified, √50 equals x√y. If x and y are both integers, what is the value of x? a)1 b)2 c)5 d)25

OpenStudy (anonymous):

\[\sqrt{50}=\sqrt{5^2\times 2}=\sqrt{5^2}\times \sqrt{2}=5\sqrt{2}\]

OpenStudy (anonymous):

so x = 5